### v1

# The Architecture and Evaluation of Bayesian Neural Networks

Abstract
As modern neural networks get more complex, specifying a model with high predictive performance and sound uncertainty quantification becomes a more challenging task. Despite some promising theoretical results on the true posterior predictive distribution of Bayesian neural networks, the properties of even the most commonly used posterior approximations are often questioned. Computational burdens and intractable posteriors expose miscalibrated Bayesian neural networks to poor accuracy and unreliable uncertainty estimates. Approximate Bayesian inference aims to replace unknown and intractable posterior distributions with some simpler but feasible distributions. The dimensions of modern deep models coupled with the lack of identifiability make Markov chain Monte Carlo tremendously expensive and unable to fully explore the multimodal posterior. On the other hand, variational inference benefits from improved computational complexity but lacks the asymptotical guarantees of sampling-based inference and tends to concentrate around a single mode. The performance of both approaches heavily depends on architectural choices; this paper aims to shed some light on this, by considering the computational costs, accuracy and uncertainty quantification in different scenarios including large width and out-of-sample data. To improve posterior exploration, different model averaging and ensembling techniques are studied, along with their benefits on predictive performance. In our experiments, variational inference overall provided better uncertainty quantification than Markov chain Monte Carlo; further, stacking and ensembles of variational approximations provided comparable to Markov chain Monte Carlo accuracy at a much-reduced cost.
Keywords: Bayesian Deep Learning, Approximate Bayesian Inference.
1 Introduction
Despite the tremendous success of deep learning in areas such as natural language processing [Touvron et al., 2023] and computer vision [Krizhevsky et al., 2017, Dosovitskiy et al., 2020], often there is no clear understanding of why a particular model performs well [Zhang et al., 2021, Szegedy et al., 2014]. Even though the universal approximation theorem guarantees that a wide enough feed-forward neural network with a single hidden layer can express any smooth function [Hornik et al., 1989], in practice, constructing a model which is not only expressive but generalizes well is challenging. In contrast, the so-called no free lunch theorem [Wolpert, 1996] dictates that there is no panacea to solve every problem, and one should be careful when designing a model appropriate to the task. Many modern machine learning models are over-parametrized and prone to overfitting, especially given the limited size of the dataset. Complex problems demand exploring bigger model spaces, and there is a danger of choosing an excessively over-parametrized model which is going to overfit and has a high variance. Additionally, conventional deep models do not offer human-understandable explanations and lack interpretability [Lipton, 2018]. By default, classical neural networks do not address the uncertainty associated with their parameters and whilst there exist proposals enabling neural networks (NNs) to provide some uncertainty estimates, they are often miscalibrated [Guo et al., 2017]. As a result, these models are typically overconfident, provide a low level of uncertainty even when data variations occur [Ovadia et al., 2019], and are easily fooled and are susceptible to adversarial attacks [Szegedy et al., 2014, Nguyen et al., 2015]. At the same time, reliable uncertainty quantification (UQ) is crucial for any decision-making process, and it is not enough to obtain a point estimate of the prediction. The key distinguishing property of the Bayesian framework is that it incorporates domain expertise and deals with uncertainty quantification in a principled way: by marginalizing over the posterior distribution of parameters. As a result, Bayesian models are more resistant to distribution shifts and can improve the accuracy and calibration of classical deep models [Wilson and Izmailov, 2020]. Nevertheless, the reliability of uncertainty estimates and the gap within-the-sample and out-of-sample performance still require improvement [Foong et al., 2020]. The distributions arising in Bayesian neural networks (BNNs) are typically infeasible and highly multimodal, and the core challenge lies in estimating the posterior [Papamarkou et al., 2022]. One should not only find a model that matches the task but, as importantly, achieve the alignment between the model and the applied inference algorithm [Gelman et al., 2020]; and the most theoretically grounded sampling methods and approximation techniques are limited by the computing budget, size of the dataset, and sheer number of parameters. We list several characteristics of classical and Bayesian neural networks in the Table 1.
Table 1: Some of the challenges and properties of classical and Bayesian neural networks.
Property Classical NN Bayesian NN
[…]
Training outcome point estimate $\cdot$ posterior distribution $\mathbb{P}$
Incorporate prior no yes
Require initialization yes yes
Outline. In this work, we consider some of the challenges and nuances of Bayesian neural networks and evaluate the performance with different architectures and for different posterior inference algorithm choices. Specifically, we study the sensitivity of BNNs to the choice of width in Section 2.3, depth in Section 2.4, and investigate the performance of BNN under the distribution shift in Section 2.5; Across all the experiments in Section 2, we observe that for different inference algorithms, one model can provide strikingly diverse performances. The challenge of comparative model assessment is addressed in Section 3.1, where we introduce the estimated pointwise loglikelihood as a measure of model utility. While, given some set of models, the Bayesian approach has the potential to deal with the model choice by comparing posterior model probabilities, such comparison tends to favour one candidate disproportionally strongly [Oelrich et al., 2020]. Thus, the classical Bayesian model averaging (BMA) based on model probabilities [Hoeting et al., 1999] is only optimal if the true model is among the comparison set. In response to the limitations of BMA, in Sections 3.2 and 3.3 we consider ensembling, stacking and pseudo-BMA[Yao et al., 2018].
2 Empirical Study of Limiting Scenarios
2.1 Architecture Components
Whilst the dimensions of the input and the output are determined by the dimensionality of the data set, the dimension of the weight space plays an essential part in specifying neural networks and can be tuned to improve prediction performance. In the case of feed-forward neural networks, this amounts to finding optimal depth and width. While the universal approximation theorem advocates for single-layer neural networks, deep neural networks gained popularity due to their expressiveness and tremendous success in real-world applications allowed by the increase in available computing power [Chatziafratis et al., 2020]. At the same time, the more parameters one has, the more nuanced the choice of the model becomes. No matter what the prediction task is, overly complex models suffer from the curse of dimensionality which causes not only poor performance but also computational problems. On a slightly different line, we recall the seminal result first obtained for neural networks with one hidden layer [Neal, 1995] and then extended to arbitrary depth [Matthews et al., 2018] which states that as the width of BNN tends to infinity, the distribution of the network’s output induced by the prior converges to the Gaussian process (GP) with a neural network kernel, also known as the neural network GP (NNGP); there is a similar correspondence relating GPs and distributions induced by the posterior [Hron et al., 2022]. When defining Bayesian neural networks, choosing a prior and understanding how properties and prior beliefs on the weight space translate to the functions is a major challenge. Even though we acknowledge the importance of prior specification, we do not empirically study different choices here. Note that generally, we require priors which are: (1) interpretable, e.g. we want to be able to specify the hyperparameters of the prior based on the task at hand; (2) have large support, i.e. prior should not concentrate around a small subset of the parameter space; (3) lead to feasible inference and favour reasonable approximations of the posterior and predictive distributions.
Finally, to specify any neural network one needs to choose the activation function which (apart from being nonlinear) is required to be differentiable. In our experiments, we consider a popular nowadays rectified linear unit function (ReLU) defined as $\max(0,x)$, which switches the negative inputs off and leaves the positive ones unchanged, as well as the sigmoid activation function defined as $\sigma(x)=\exp(x)/(\exp(x)+1)$.
2.2 Settings of the Experiment
In the experiment, we consider the following Bayesian neural network, illustrated by the Figure 1
Figure 1: Example of the directed acyclic graph (DAG) of the neural network used in the experiments when $L=2$.
$\bm{y}\sim\text{N}\left(\bm{b}_{L+1}+\bm{W}_{L+1}\bm{z}_{L},\bm{\sigma}\right),\quad\bm{\sigma}\sim|\text{N}(0,0.001)|$ (1)
$\bm{z}_{l}=g\left(\bm{b}_{l}+\bm{W}_{l}\bm{z}_{l-1}\right)\text{ for }l=1,\ldots,L,$
where we consider two different choices of nonlinearity $g$, namely, the ReLU and the sigmoid and the following priors on the weights and biases:
$\bm{W}_{l}\sim\text{N}(0,\frac{1}{\sqrt{LD_{0}}})\quad\bm{W}_{l}\sim\text{N}(0,\frac{2}{\sqrt{D_{l-1}}}),\text{ for }l=2,\ldots,L+1,$
$\bm{b}_{l}\sim\text{N}(0,\frac{1}{2\sqrt{L}})\text{ for }l=1,\ldots,L+1,$
[…]
2.3 Limiting the Width of the Network
We consider a simple synthetic dataset with one-dimensional input and output $\sin(10\bm{x})\bm{x}^{2}$, where the noisy observations on which the neural network is trained are obtained by adding some Gaussian noise:
$\bm{x}$ $\sim\text{Unif}([0,2]),\;\bm{y}=\sin(10\bm{x})\bm{x}^{2}+\bm{\epsilon},\;\bm{\epsilon}\sim\text{N}(0,0.25).$
The input is scaled, unlike the output. The training data $\bm{\mathcal{D}}$ consists of $N=500$ observations and the new data for testing $\tilde{\bm{\mathcal{D}}}$ consists of $\tilde{N}=100$ observations. We first study the performances of mfVIR, mfVIS, HMCR and HMCS with 1 hidden layer as the width increases and illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units by the Figure 2(a). The predictions of all four models when $D_{1}=2000$ are provided on the Figure 2(b). The performance of the mfVIS dips with the increase in the dimension of the hidden layer, moreover, for $D_{1}=1000$ and $D_{1}=2000$ its posterior predictive distribution fails to capture the data, and, in fact, degenerates to the prior (Figure 2(b)). An explanation of why such behaviour occurs was obtained via the correspondence of Gaussian processes and BNNs. While as the width increases the true posterior of a BNN converges to NNGP posterior [Hron et al., 2022], any optimal mean-field Gaussian variational posterior of a BNN with odd (up to a constant offset) Lipschitz activation function converges to the prior predictive distribution of the NNGP [Coker et al., 2022]. In other words, the mean-field variational approximations of wide BNNs with sigmoid activation function ignore the data. If one abandons the mean-field assumption and proposes a full-rank variational family, then using variational inference (VI) for wider networks would take at least a hundred times more time than using HMC, which undermines the benefits of using VI. Such degenerate behaviour is not observed with HMC,(Figure 2(b)), but this comes at a subsequent significant increase in training time. For wider networks, the HMCR model exhibits a better performance than the HMCS both in terms of accuracy and uncertainty quantification. In terms of predictive accuracy, HMC is preferred over mfVI in all of the combinations of the activation function and width. However, in terms of uncertainty quantification, the HMC is inferior to mfVI. In our experiment, HMC underestimates the uncertainty of the signal much more than VI (Figures 2(b) and 2(a)). Note that whilst variational inference is often cursed to underestimate the uncertainty[Trippe and Turner, 2018], that is not always the case [Blundell et al., 2015, Gal and Ghahramani, 2016]. Markov chain Monte Carlo (MCMC) methods are known to struggle to effectively explore multimodal posteriors [Papamarkou et al., 2022, Izmailov et al., 2021], and lack of uncertainty could be a result of poor mixing of the chain.
General summary. In wider networks, the ReLU is preferred over the sigmoid activation for both HMC and mfVI. Crucially, when it comes to the mean-field VI the sigmoid activation should only be used when the limited width is suitable for the task at hand. It is reasonable to suppose that the same could be said about any odd (up to adding a constant) activation function. Further, while the HMC was preferred over the mfVI when looking at accuracy alone, the required computational resources could be an obstacle. Moreover, uncertainty quantification is far from ideal for HMC (CIs are too narrow for the signal); instead, mfVI with the ReLU achieves a good balance between accuracy, UQ, and time, particularly for wider networks.
(a) Performance of models as the number of hidden units increases.
(b) The predictions and uncertainty estimates obtained by each model when $D_{1}=2000$
Figure 2: Predictive performance of wider BNNs.
2.4 Deeper Networks
Consider the data of Section 2.3 and neural networks defined by Equation 1 with the number of layers $L$ varying from 1 to 6 and a fixed number of hidden units in each layer $D_{h}=20$. Figure 3(a) provides the recorded metrics, and Figure 3(b) illustrates the predictions of all the four models with $L=6$. First, observe that overall both RMSE and empirical coverage of mfVIR approximations improve with the increase of depth. The mfVIS follows a similar pattern, except for the $L=5$ when the prediction quality of the network drops drastically. Indeed, the approximate posteriors of deep neural networks obtained with the mean-field variational inference were shown to be as flexible as the much richer approximate posteriors of shallower BNNs [Farquhar et al., 2020]. We do not obtain the same improvement in the prediction quality of models trained with HMC: the performance of HMCR falls whilst the HMCS does not improve as the depth increases. This undesirable behaviour could be a result of the multimodality of distributions in overparametrized models combined with the challenges of MCMC in exploring the high-dimensional space[Izmailov et al., 2021, Papamarkou et al., 2022]. Compared to the findings of Section 2.3, we note that the deeper NNs are less sensitive to the choice of the activation function. It is needless to say that the HMC algorithm scales rather poorly and as the number of layers changes from $L=1$ to $L=6$ the time needed to train HMCR and HMCS gets more than 15 and 30 times greater, respectively. We note that for models with more than one hidden layer, the network with sigmoid activations takes roughly twice as much time as the network with ReLU. The striking discrepancy in training times could arise due to the difference in the leapfrog integrator step sizes [Betancourt et al., 2015].
(a) Performance of models as the number of hidden layers increases.
(b) The predictions and uncertainty estimates obtained by each model when $L=6$
Figure 3: Prediction performance of deeper networks.
General summary. In terms of the training time, HMC becomes less and less feasible with the increase in depth. With the need to explore high-dimensional parameter spaces, multimodality of the posteriors should be kept in mind as an arising challenge for both mfVI and HMC. In terms of the balance between accuracy and UQ, the mean-field variational inference with ReLU activation function is able to outperform the MCMC with the increase in depth.
2.5 Out-of-Distribution Prediction
While it is not surprising that the accuracy and the quality of uncertainty quantification of any model decreases under a distribution shift, reliable uncertainty estimates that are robust to the out-of-distribution (OOD) data become exceptionally important in safety-critical applications. The challenge is especially intricate since better accuracy and lower calibration error of a certain model on the in-domain data do not imply better accuracy and lower calibration error in the OOD settings [Ovadia et al., 2019]. Here, we wish to validate the models’ predictive abilities when the test data points come from previously unseen regions of data space. The kind of out-of-distribution data we consider could be described as ’complement-distributions’, such data arises in open-set recognition or could be the result of adversary [Farquhar and Gal, 2022]. Note that in Section C.1 we introduce a much milder example with ’related-distributions’ test data. We split the training data used in Sections 2.3 and 2.4 into the train and test data covering complement regions of the function. Specifically, $\bm{\mathcal{D}}=\bm{\mathcal{D}}_{c}\sqcup\tilde{\bm{\mathcal{D}}}_{c}$, the observed data $\bm{\mathcal{D}}_{c}$ consists of $N=370$, the new data $\tilde{\bm{\mathcal{D}}}_{c}$ consists of $\tilde{N}=130$ and the observed and the new data are disjoint (see Figure 4):
$\bm{\mathcal{D}}_{c}$ =$=\{(x_{n},y_{n})\;|\;x_{n}\in[-1.7,1.7]\},$
$\tilde{\bm{\mathcal{D}}}_{c}$ =$=\{(x_{n},y_{n})\;|\;x_{n}\in[-2.8,-1.7)\cup(1.7,1.9)]\}.$
Strictly speaking, we do not expect any model to be robust to such an extreme case and, mainly, want to asses and better understand the quality of the uncertainty estimates. In this experiment, we are hoping that the relationship between the distributions of the observed and the new data makes this challenge somewhat tractable. On Figure 4(a) we illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units; Figure 4(b) compares non-OOD and OOD predictions obtained by the BNNs with ReLU activation and $D_{1}=200$. The poor performance of the mfVIS, especially for wider networks is not surprising, however, we notice that HMCS suffers from much higher RMSE than mfVIR and HMCR for wide networks. And while HMCR has a lower RMSE than any model trained with mean-field VI, the ability of HMC to capture the uncertainty deteriorates and it becomes overconfident. Whilst HMCR200 and mfVIR200 do not show any of the expected increase in the uncertainty, on certain regions both methods are able to provide accurate predictive mean (see Figure 4(b), the RHS region of the function, where $x>1.5$). Finally, as the width of the network increases, mfVIR outperforms all of the models.
(a) Out-of-sample predictive performance of all the models compared.
(b) Within-the sample and out-of-sample predictions and uncertainty estimates of BNNs with $D_{1}=200$ and ReLU activation.
Figure 4: Out-of-distribution prediction for the ’complement-distributions’ data.
General summary. In terms of the accuracy alone, the HMC with ReLU is more robust to the out-of-distribution data, however, that comes with the largest computational costs among all the models. We already saw in Section 2.3 that uncertainty quantification with HMC degrades with increasing width. In OOD settings, this becomes even more extreme, with very overconfident predictions that do not cover the truth (an empirical coverage of almost zero). Finally, with the increase in depth, in the extreme OOD settings, the mfVI with ReLU becomes almost as accurate as HMC with ReLU and provides better UQ at a much lower cost.
3 Bayesian Model Averaging and Stacking
3.1 Predictive Methods for Model Assessment
When considering synthetic datasets, we can choose a desired metric and sample any number of data points, so that evaluation of the model’s performance becomes trivial. For example, in Section 2.5 we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. In reality, the new previously unseen data is not available, and one can only estimate the expected out-of-sample predictive performance. Suppose that we only observe $\bm{\mathcal{D}}$, the unseen observations $\tilde{\bm{\mathcal{D}}}$ are generated by $p_{t}(\tilde{\bm{\mathcal{D}}})$, and we wish to be able to access the generalization ability of the model without having access to the test data. To keep the notation simple we omit the dependency on $\bm{x}$ and $\tilde{\bm{x}}$ when writing down the posteriors in this section. Given a new data point $\tilde{y}_{n}$, the log score $\log p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is one of the most common utility functions used in measuring the quality of the predictive distribution. The log score benefits from being a local and proper scoring rule [Vehtari and Ojanen, 2012]. Then, the expected log pointwise predictive density for a new dataset serves as a measure of the predictive accuracy of a given model:
$\text{elpd}=\sum_{n=1}^{\tilde{N}}\int p_{t}(\tilde{\mathrm{D}}_{n})\log p(\tilde{y}_{n}|\bm{\mathcal{D}})\text{d}\tilde{\mathrm{D}}_{n},$
where $p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is model’s posterior predictive distribution. In the absence of $\tilde{\bm{\mathcal{D}}}$, one might obtain an estimate of the expected log pointwise predictive density by re-using the observed $\bm{\mathcal{D}}$. Here, we review the approach that employs leave-one-out cross-validation (LOO-CV), which can be seen as a natural framework for accessing the model’s predictive performance [Vehtari et al., 2016].
Willing to obtain the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility $\widehat{\text{elpd}}_{\text{loo}}$ and avoid re-fitting the model $N$ times one could use importance sampling. However, the classical importance weights would have a large variance and the obtained estimates would be noisy. Recently, the problem was solved with the Pareto smoothed importance sampling (PSIS) which allows evaluating the LOO-CV expected utility in a reliable yet efficient way [Vehtari et al., 2016]:
$\widehat{\text{elpd}}_{\text{loo}}$ $=\sum_{n=1}^{N}p(y_{n}|x_{n},\bm{\mathcal{D}}_{-n})=\sum_{n}^{N}\log\left(\frac{\sum_{s=1}^{S}r_{i}^{s}p(y_{n}|\bm{\theta}^{s})}{\sum_{s=1}^{S}r_{i}^{s}}\right),$ (2)
where $r_{i}^{s}$ are the smoothed importance weights which benefit from smaller variance than the classical weights. We refer to the individual logarithms in the sum as to $\widehat{\text{elpd}}_{\text{loo},n}$. The advantage of PSIS is that the estimated shape parameter of the Pareto distribution provides a diagnostic of the reliability of the resulting expected utility. Although the methods of model selection which reuse the data can be vulnerable to overfitting when the size of the dataset is too small and/or the data is sparse, it is (relatively) safe to use cross-validation to compare a small number of models and given a large enough dataset [Vetari et al., 2019]. In Section B.1, we implement $\widehat{\text{elpd}}_{\text{loo}}$ in the empirical experiment, where we additionally consider posterior predictive checks (PPC) and an alternative to the LOO-CV approach of estimating the expected log pointwise utility.
3.2 Alternatives to Classical Bayesian Model Averaging
Let =$\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ be a collection of models and denote the parameters of each of the $M_{k}$ as $\bm{\theta}_{k}$. The assumptions one has on the prediction task and on $\mathcal{M}$ with respect to the true data-generating process, can be categorized into three scenarios: $\mathcal{M}$-closed, $\mathcal{M}$-open and $\mathcal{M}$-complete. If $M_{k}\in\mathcal{M}$ for some $k$ recovers the true data generating process then we are in the $\mathcal{M}$-closed case. The task is $\mathcal{M}$-complete if there exists a true model but it is not included in $\mathcal{M}$ (e.g. for computational reasons). Finally, we are in the $\mathcal{M}$-open scenario when the true model is not in $\mathcal{M}$ and the data generating mechanism cannot be conceptually formalized to provide an explicit model [Vehtari and Ojanen, 2012]. The Bayesian framework allows to define the probabilities over the model space and the $\mathcal{M}$-closed case, the classical Bayesian Model Averaging would be able to give optimal performance. The BMA solution provides an averaged predictive posterior as [Hoeting et al., 1999]
$p(\tilde{\bm{y}}\mid\bm{\mathcal{D}})$ $=\sum_{k=1}^{K}p(\tilde{\bm{y}}\mid\bm{\mathcal{D}},M_{k})p(M_{k}\mid\bm{\mathcal{D}}),$ (3)
where $p(M_{k}\mid\bm{\mathcal{D}})\propto p(\bm{\mathcal{D}}\mid M_{k})p(M_{k}).$ (4)
However, in the $\mathcal{M}$-open and $\mathcal{M}$-complete prediction tasks, the BMA is not appropriate as it gives a strong preference to a single model and so assumes that this particular model is the true one. Now, if we replace the weights $p(M_{k}|\bm{\mathcal{D}})$ with the products of Bayesian LOO-CV densities $\prod_{n=1}^{N}p(y_{n}\mid x_{n},\bm{\mathcal{D}}_{-n},M_{k})$, we arrive at pseudo-Bayesian model averaging (pseudo-BMA). In other words, the weights $w_{k}$ of pseudo-BMA are proportional to the estimated log pointwise predictive density $\exp(\widehat{\text{elpd}}_{\text{loo}}^{k})$ introduced in Section 3.1. One could further correct each $\widehat{\text{elpd}}_{\text{loo}}$ estimate of Equation 2 by the standard errors and obtain
$w_{k}$ $=\frac{\exp(\widehat{\text{elpd}}_{\text{loo}}^{k,\text{reg}})}{\sum_{k=1}^{K}\exp(\widehat{\text{elpd}}_{\text{loo}}^{k,\text{reg}})},$
$\widehat{\text{elpd}}_{\text{loo}}^{k,\text{reg}}$ $=\widehat{\text{elpd}}_{\text{loo}}^{k}-\frac{1}{2}\sqrt{\sum_{n=1}^{N}\left(\widehat{\text{elpd}}_{\text{loo},n}^{k}-\frac{\widehat{\text{elpd}}_{\text{loo}}^{k}}{n}\right)^{2}},$
where for each model $M_{k}$ we find $\widehat{\text{elpd}}_{\text{loo}}^{k,\text{reg}}$ by utilizing a log-normal approximation. Fortunately, we have already seen that these densities can be efficiently estimated with PSIS.
An alternative way to obtain the averaged predictive posterior given the set of $p(\tilde{\bm{y}}\mid\bm{\mathcal{D}},M_{k})$ is to employ the stacking approach [Yao et al., 2018]. Define the set $S^{K}=\{\bm{w}\in[0,1]^{K}|\sum_{k}^{K}w_{k}=1\}$, then the stacking weights are found as the optimal (according to the logarithmic score) solution of the following problem
$\bm{w}$ $=\max_{\bm{w}\in S^{K}}\frac{1}{N}\sum_{n=1}^{N}\log\sum_{k=1}^{K}w_{k}p(y_{n}\mid\bm{\mathcal{D}}_{-n},M_{k}),$
$=\max_{\bm{w}\in S^{K}}\frac{1}{N}\sum_{n=1}^{N}\log\sum_{k=1}^{K}w_{k}\left(\frac{\sum_{s=1}^{S}r_{i}^{s}p(y_{n}\mid\bm{\theta}_{k}^{s},M_{k})}{\sum_{s=1}^{S}r_{i}^{s}}\right),$
where a PSIS estimate of the predictive LOO-CV density is used and $r_{i}^{s}$ are the smoothed (truncated) importance weights.
Finally, we recall that deep ensembles of classical non-Bayesian NNs [Lakshminarayanan et al., 2017] behave similarly to Bayesian model averages, and both lead to solutions strongly favouring one single model [Wilson and Izmailov, 2020]. In contrast, the ensembles of BNN posteriors in the Equation 3 with $p(M_{k}\mid\bm{\mathcal{D}})=K^{-1}$ can be seen as a trivial case of BMA which combines models and does not give preference to a single solution. Alternatively, when implementing variational inference and combining BNNs, the analogy can be drawn with the simplified version of the adaptive variational Bayes, which combines variational posteriors with certain weights and under certain conditions, attains optimal contraction rates [Ohn and Lin, 2024].
3.3 Ensembles and Averages
We compare three model averaging methodologies: deep ensembles of Bayesian neural networks, stacking and pseudo-BMA based on PSIS-LOO [Yao et al., 2018]. We do not consider the Bayesian Bootstrap (BB) [Rubin, 1981] motivated by the recent observation that in the settings of modern neural networks deep ensembles of non-Bayesian NNs and BB are equivalent, and both are often misspecified [Wu and A Williamson, 2024]. Combining several estimates of BNNs can be effective not only when predictions are coming from different models, but also when dealing with several predictions obtained by the same model [Ohn and Lin, 2024]. This is of particular use for multimodal posteriors arising in BNNs where different modes could be explored by random initializations [Yao et al., 2018]. Additionally recall, that the ELBO, the objective of variational inference, is a non-convex function so that the optimum is only local and depends on the starting point. We note that combining models trained with HMC and VI would be meaningless for several reasons. First of all, training a set of HMC models becomes rather expensive: for instance, training the HMCR20 once takes the same amount of time as 35 trainings of mfVIR20. Second, the estimates of the log pointwise predictive densities (provided in Section B.1) for HMC and VI have different scales and are not easily compared, in this case, the result of averaging HMC and VI would be equivalent to classical BMA. The reader, nevertheless, interested in ensembles and averages of HMC runs is referred to Section C.2 where we show that in this particular example, neither of the methods promoted exploring the multiple modes of the posterior predictive distribution nor did they help improve uncertainty quantification. Additionally Section C.3 provides results of ensembling and averaging the deeper networks and Section C.1 tests the performance of averaging methods in less extreme examples of OOD data by introducing a much milder example with the so-called ’related-distributions’ test data Farquhar and Gal [2022].
Now consider the mfVIR20 model and the ’complement-distributions’ data of Section 2.5. We choose 10 random initialization points, obtain 10 posterior predictive distributions and compute estimated expected log pointwise predictive densities. We then construct ensemble, pseudo-BMA and stacking approximations, the results are illustrated Figure 5). The ensembling and stacking provide subtle results and are superior to pseudo-BMA which has worse accuracy and fails to capture any uncertainty..
Figure 5: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20. The pseudo-BMA is worse than DE and stacking which are very similar to each other.
General summary. We observe that, similarly to BMA, the pseudo-BMA is not preferable in $\mathcal{M}$-open and $\mathcal{M}$-complete settings. Namely, in the experiments the pseudo-BMA was confirmed to be inferior to stacking and ensembles of BNNs both, in terms of the predictive accuracy and the uncertainty quantification. Stacking and ensembles of BNNs performed comparable to each other and provided an improvement, which is especially significant in terms of the uncertainty quantification.
4 Discussion
The message of an optimist’s conclusion could question the common belief that the mean-field variational approximations are generally overly restrictive and do not capture the true posterior and the uncertainty well. Even given the increased available computing power and assuming that HMC can be taken as a gold standard (which as we have seen is not the case), computational costs of the sampling algorithms suggest that it may not be feasible for most of the modern neural networks and datasets. Indeed, in a variety of the experiments considered in Sections 2.3, 2.4, 2.5 and 3.3 mfVI overall provided better uncertainty quantification than HMC, and in out-of-distribution setting the empirical coverage of the latter was close to zero. We note that for single-layer neural networks HMC outperformed mfVI only in terms of accuracy. At the same time, for deeper networks and in out-of-distribution scenarios, the accuracy of mfVI was often comparable to HMC. Further, in Section 2.4 we confirmed that even for slightly deeper networks the time needed to perform HMC becomes a burden which makes variational inference a very attractive alternative to sampling. Nevertheless, in Section 2.3 we observed that the restrictions imposed by the factorized families can obstruct models from effectively learning from the data. In real-life scenarios where one is required to evaluate the future predictive performance of the model before applying it to the unseen data, the estimate of the expected log pointwise predictive density can serve as a reliable diagnostic and thus, PSIS-LOO estimates can be beneficial for model assessment and combination. In Section 3.3, stacking as well as, ensembles of BNNs were shown to be a possible solution when dealing with multimodal posteriors, helping to both improve accuracy and uncertainty quantification even in the extreme OOD scenario. We find that stacked or ensembled variational approximations are competitive to HMC at a much-reduced cost.
This work highlights the model’s sensitivity to architectural choices, namely, width, depth and activation function. Along the same line, future work could study the performance of various priors, including sparsity-inducing priors which are known to reduce computational costs whilst improving the accuracy and calibration [Blundell et al., 2015, Polson and Ročková, 2018]. Further, an important avenue for research is to consider the so-called structured variational inference with less restrictive variational families, and more generally, study the trade-off between the expressiveness of the variational family and scalability. Finally, given the multimodal nature of distributions arising in Bayesian neural networks, a promising avenue for research is to continue improving model combination techniques, possibly along the lines of an adaptive variational Bayes framework [Ohn and Lin, 2024] or hierarchical stacking and combining the models pointwise [Yao et al., 2022].
Appendix A Metrics and Practicalities
Recall that we denoted the training data to be $\bm{\mathcal{D}}=\{x_{n},y_{n}\}_{i=1}^{N}$ and the new data for testing to be $\tilde{\bm{\mathcal{D}}}=\{\tilde{x}_{n},\tilde{y}_{n},\}_{n=1}^{\tilde{N}}$. Denote the approximated posterior predictive mean as $\bm{y}$, the set of $S$ samples of the signal as $\bm{\mu}^{S}$ and of the observations as $\bm{y}^{S}$. Upon computing the posterior predictive distribution, we obtain the root mean squared error and the empirical coverage as:
RMSE $=\sqrt{\frac{1}{N}\sum_{n}^{N}\left[(\tilde{y}_{n}-\mathbb{E}^{S}[y^{S}_{i}])^{2}\right]},$
EC =#$=\frac{\#\{\bm{y}_{\in}[q_{0.025},q_{0.975}]\}}{N},\text{ where }q\text{ are quantiles of }\bm{\mu}^{S}\text{ or }\bm{y}^{S}.$
The results are obtained when the number of iterations of mfVI is set $10^{4}$, $10^{4}$, $5\times 10^{4}$, $6\times 10^{4}$ to train models with, respectively, $20,200,1000,2000$ hidden units in a layer. In models trained with HMC, the number of samples used for warmup was set to $10^{3}$, the samples used for posterior is $10^{3}$ in models with $20$ hidden units, and $2\times 10^{3}$ in models with $200,1000,2000$ hidden units in a layer. In experiments the number of iterations of mfVI was set to $L\times 10^{4}$; the HMC had the number of warmup samples fixed to $10^{3}$ and the number of samples was $\min(4\times 10^{3},L\times 10^{3})$.
Remark on the initialization. Based on the empirical evidence, we observed that in our experiment for $L=1,2$ the NumPyro implementation of mfVI requires the initialization mode to be set to "init to feasible", which chooses the initialization point uniformly (ignoring the prior distribution). Whereas for $L=3,4,5,6$ mfVI requires "init to mean", which sets initial parameters to the prior mean, and "init to feasible" will fail. Conversely, the NumPyro implementation of the HMC fails if the initialization location is set to "init to mean" but performs fine if it is always set to "init to feasible", i.e. ignoring the distribution parameters.
Appendix B Supplementary to Predictive Model Assessment
B.1 Empirical model assessment
Recall, the OOD scenario, where we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. To assess the expected out-of-sample predictive performance, we could begin with posterior predictive checks (PPC), which compare the true $\bm{\mathcal{D}}_{c}$ to datasets simulated from the posterior predictive distribution Gelman et al. [2020]. Figure 6(a) provides the PPC based on the kernel density estimates of the observed $\bm{y}$ and Figure 6(b) compares posterior predictive to the observed $\bm{y}$ for all of the models with $D_{1}=2000$; both figures evidently classify the mean-field VI with sigmoid activation as inappropriate. However, it is not apparent that HMCS2000 is significantly inferior to HMCR2000. We compare the RMSE computed in the OOD settings to $\widehat{\text{elpd}}_{\text{loo}}$ estimates. Since the idea of estimating the expected log predictive density is in evaluating future predictive performance, we expect the higher $\widehat{\text{elpd}}_{\text{loo}}$ to correspond to a better model and lower RMSE. Indeed, Figure 6(c) is more informative in this sense than PPC diagnostics, and we observe the inverse dependence between $\widehat{\text{elpd}}_{\text{loo}}$ and RMSE. Sampling and approximation techniques result in different scales of $\widehat{\text{elpd}}_{\text{loo}}$ estimates (in general, $\widehat{\text{elpd}}_{\text{loo}}$ is lower for VI, especially, in wide networks) and thus, we compare the models trained with different algorithms for better visualization.
(a) Posterior predictive checks for the wider models based on the kernel density estimates of the observed $\bm{y}$.
(b) Posterior predictive distribution and posterior predictive mean compared to the observed $\bm{y}$.
(c) The correspondence between the $\widehat{\text{elpd}}_{\text{loo}}$ and the RMSE in the OOD scenario. Higher $\widehat{\text{elpd}}_{\text{loo}}$ should correspond to lower RMSE.
Figure 6: Estimating the out-of-distribution performance before seeing the new data: testing the (a), (b) PPC and (c) $\widehat{\text{elpd}}_{\text{loo}}$. The mfVIR2000 is confirmed to be unreliable in all methods. The PPC of the HMCS2000 does not provide enough information to judge its performance in the OOD settings, while the $\widehat{\text{elpd}}_{\text{loo}}$ does.
General summary. In certain cases, such as mfVIS for large width, the posterior predictive checks are able to detect an undesirable model. However, when the PPCs are not sufficient, we confirmed that the PSIS-LOO estimates of the expected log pointwise predictive density can serve as robust diagnostics for both the mfVI and the HMC methods.
B.2 Correspondence Between WAIC and RMSE
Above, we compared the RMSE obtained in the OOD experiment to the estimates of the expected log pointwise predictive density obtained with LOO-CV. An alternative approach is to overestimate the elpd by first computing the log pointwise predictive density of $\bm{\mathcal{D}}$ and then adjusting it by some correction term. Specifically, suppose that $\bm{\theta}$ are the parameters of the model and $\bm{\theta}^{s},s=1,\ldots,S$ are simulation draws. Then, one can evaluate the log pointwise predictive density as
$\widehat{\text{lpd}}$ $=\sum_{n=1}^{N}\log\left[\frac{1}{S}\sum_{s=1}^{S}p(y_{n}|\bm{\theta}^{s})\right].$
With $\widehat{\text{lpd}}$ at hand, a superior successor of the Akaike Information criterion (AIC) Akaike [1998] and the Deviance information criterion (DIC) Spiegelhalter et al. [2002] called the Watanabe-Akaike information criterion (WAIC)  [^2] Also called Widely Applicable information criterion Watanabe [2010] can be obtained. To mitigate the bias between the log pointwise density and the expected utility, WAIC subtracts the simulation-estimated effective number of parameters:
$\widehat{\text{elpd}}_{\text{WAIC}}=\widehat{\text{lpd}}-\widehat{p}_{\text{WAIC}},\text{ where }\widehat{p}_{\text{WAIC}}=\sum_{n=1}^{N}\operatorname{Var}^{S}(p(y_{n}|\bm{\theta}^{s})).$
Here, $\operatorname{Var}^{S}$ is the sample variance and the estimated effective number of parameters $\widehat{p}_{\text{WAIC}}$ can be seen as a measure of model complexity. Asymptotically, WAIC is equivalent to the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility Watanabe [2010]. Even though cross-validation is a natural framework for accessing the model’s predictive performance, the WAIC was for a long time preferred over the LOO-CV due to the computational challenges arising from multiple model runs Gelman et al. [2014]. Moreover, whilst both the PSIS-LOO and WAIC estimates give nearly unbiased estimates of the predictive ability of the model, $\widehat{\text{elpd}}_{\text{loo}}$ was shown to be more robust than $\widehat{\text{elpd}}_{\text{WAIC}}$; in the presence of limited sample size and weak priors, WAIC can severely underestimate $\hat{p}_{\text{WAIC}}$ and often has a larger bias towards the log predictive density Vehtari et al. [2016], Gelman et al. [2014].
That said, computing $\widehat{\text{elpd}}_{\text{loo}}$ involved Pareto smoothed importance sampling, and in some of the models, the estimated shape parameter of the generalized Pareto distribution gave a warning about the reliability of the LOO estimate. Here we do an extra step and check if the WAIC and LOO estimates of the elpd agree. Figure 7 illustrates the reverse dependency between the RMSE and $\widehat{\text{elpd}}_{\text{WAIC}}$ and is largely identical to the figure illustrating the dependency between the RMSE and $\widehat{\text{elpd}}_{\text{loo}}$ (in terms of the location of coordinates but not the error bars). Therefore, we can conclude that LOO estimates of the elphd can be seen as relatively reliable.
Figure 7: Estimating the out-of-distribution performance before seeing the new data: the correspondence between the $\widehat{\text{elpd}}_{\text{WAIC}}$ and the RMSE in the OOD scenario. Similarly to $\widehat{\text{elpd}}_{\text{loo}}$, the higher $\widehat{\text{elpd}}_{\text{WAIC}}$ should correspond to lower RMSE.
Appendix C Supplementary to Ensembles and Averages
Remark on constructing deep ensembles. Given =$\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ a collection of models suppose that $K$ approximations $\tilde{\bm{y}}_{k}$ of the posterior $p(\tilde{\bm{y}}|\bm{\mathcal{D}},M_{k})$ have means $\mu_{k}$ and variances $\sigma_{k}^{2}$ or $k=1,\ldots,K$. Then the mean and variance of an ensemble of approximations:
$\mu_{\text{DE}}=K^{-1}\sum_{1}^{K}\mu_{k},\quad\sigma_{\text{DE}}^{2}=K^{-1}\sum_{1}^{K}\sigma_{k}^{2}+\mu_{k}^{2}-\mu_{\text{DE}}^{2}.$
In general, given weights $\omega_{k}=p(M=M_{k})$ the mean and the variance are
$\mu_{\text{DE}}$ $=\mathbb{E}[\mathbb{E}[\tilde{\bm{y}}|M]]=\sum_{1}^{K}\omega_{k}\mu_{k},$
$\sigma_{\text{DE}}^{2}$ $=\mathbb{E}[\mathbb{E}[\tilde{\bm{y}}^{2}|M]]-\mu_{\text{DE}}^{2}=\sum_{1}^{K}\omega_{k}\mathbb{E}[\tilde{\bm{y}_{k}}^{2}]-\mu_{\text{DE}}^{2}$
$=\sum_{1}^{K}\omega_{k}(\sigma_{k}^{2}+\mu_{k}^{2})-\mu_{\text{DE}}^{2}.$
C.1 Milder OOD Scenario
Given the nature of the test data we use, it is reasonable to assume that not only the predictions but also the $\widehat{\text{elpd}}_{\text{loo}}$ estimates are tentative. Thus, in order to explore different methodologies from different angles we create a simpler regression problem in which test data comes from a slightly broader region. Using the classification of Farquhar and Gal [2022], this new scenario could be named as an OOD task with ’related-distributions’. We define a similar synthetic dataset with one-dimensional input and output, to which we add some small noise:
$\bm{x}$ $\sim\text{Unif}([0,1])$
$\bm{y}$ $=\sin(10\bm{x})\bm{x}^{2}+\bm{\epsilon}$
$\bm{\epsilon}$ $\sim 0.05\text{N}(0,1)$
As before, the input is scaled, unlike the output. The data for training $\bm{\mathcal{D}}_{r}$ and the testing $\tilde{\bm{\mathcal{D}}}_{r}$ consist of $N=450$ and $\tilde{N}=50$ observations, respectively, where $\tilde{\bm{\mathcal{D}}}_{r}$ comes from the broader than $\bm{\mathcal{D}}_{r}$ region, i.e. $(\min_{n=1\ldots N}(x_{n}),\max_{n=1\ldots N}(x_{n}))\subset(\min_{n=1\ldots\tilde{N}}(\tilde{x}_{n}),\max_{n=1\ldots\tilde{N}}(\tilde{x}_{n}))$. Having 10 posterior predictive distributions of mfVIR20, we compare ensembling, pseudo-BMA and stacking on Figure 8. Whilst the total uncertainty estimates of pseudo-BMA are, somewhat, adequate, the model uncertainty is underestimated. Stacking provides slightly better accuracy and empirical coverage than the deep ensembles, with both techniques leading to improved prediction performance and uncertainty quantification.
Figure 8: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 in the ’related-distributions’ task. The pseudo-BMA is again worse than the other methodologies. Stacking is slightly better than DE.
C.2 HMC and Averaging Techniques
We wish to recreate the experiment for the HMCR20 model (instead of the mfVIR20 mode). We choose 10 random initialization points, obtain 10 posterior predictive distributions and compute estimated expected log pointwise predictive densities. We then construct ensemble, pseudo-BMA and stacking approximations, for the ’complement-distributions’ task the results are illustrated by Figure 9 and the predictions of the ’related-distributions’ task are shown on Figure 10. In contrast to mfVIR20, this time we do not observe a clear difference between the pseudo-BMA and stacking and ensembling methodologies. Moreover, all the approaches require considerable computational resources (for 10 random runs) but do not provide a considerable improvement in RMSE and empirical coverage compared to a single random run of the model. We conclude that in this particular example, ensembling, stacking and pseudo-BMA do not help explore different modes of the posterior and so cannot be recommended when dealing with HMC.
Figure 9: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to HMCR20 in the ’complement-distributions’ task are very similar.
Figure 10: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to HMCR20 in the ’related-distributions’ task are (again) strikingly similar.
C.3 Deeper Networks
Based on the 10 posterior predictive distributions obtained starting from 10 different random initialization points, we construct ensemble, pseudo-BMA and stacking approximations for mfVIR and mfVIS models with $L=6$ hidden layers. The results are consistent with the observation made in the main body of the work; in the ’complement-distributions’ task (Figure 12) pseudo-BMA is confirmed to be inferior to stacking and deep ensembles of BNNs. In the related distribution task (Figure 11), in terms of both accuracy and uncertainty quantification, stacking is preferable over deep ensembles and pseudo-BMA with the latter performing better than ensembles (unlike in the one-layer case).
Figure 11: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with $L=6$ in the ’complement-distributions’ task.
Figure 12: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with $L=6$ in the ’related-distributions’ task.